Determine the value c so that each of the following functions can serve as a probability distribution of the discrete random variable X: (b) f(x) = c (2 ) (3 )

( assume 2 & x in same parentheses and 3&3-x) ( x ) (3 - x) , for x = 0, 1, 2


1
Expert's answer
2021-09-21T17:46:52-0400

Given the function "f(x)=c*(x)*(3-x)" and "x=0,1,2" ,values for the function of these values are determined by substituting the "x's" in the function "f(x)" as shown below,

"f(0)=c*(0)*(3-0)=0, \nf(1)=c*(1)*(3-1)=2c,\nf(2)=c*(2)(3-2)=2c."

For it to be a probability distribution, it must satisfy the condition, "summation(f(x))=1, \forall x"

so,

0+2c+2c=1

4c=1

c=1/4

Therefore, a constant c=1/4, makes this function a probability distribution.